OpenStudy (anonymous):

Solve by method of GCF: 2x+5x^2=0

OpenStudy (cwrw238):

gcf is greatest common factor of the 2 terms 2x and 5x^2 can you find this?

OpenStudy (cwrw238):

2 and 5 have no common factor but is there a gcf for x and x^2?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

But I don't know what to do with the x

OpenStudy (cwrw238):

right so x goes outside of brackets and equation becomes x(2 + 5x) = 0 see how i got that?

OpenStudy (anonymous):

yeah that makes sense so far

OpenStudy (cwrw238):

if you expand the brackets you get back to original equation

OpenStudy (cwrw238):

when 2 functions when multiplied together = 0 then one or the other must be 0

OpenStudy (cwrw238):

so either x = 0 or 2 + 5x = 0 can u finish this off?

OpenStudy (anonymous):

so x=-2/5 or x=0?

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

ohmygod. Thank you so much

OpenStudy (cwrw238):

yw